### v1

@@ Abstract
Many incurable diseases prevalent across global societies stem from various influences, including lifestyle choices, economic conditions, social factors, and genetics. Research predominantly focuses on these diseases due to their widespread nature, aiming to decrease mortality, enhance treatment options, and improve healthcare standards. Among these, kidney disease stands out as a particularly severe condition affecting men and women worldwide. Nonetheless, there is a pressing need for continued research into innovative, early diagnostic methods to develop more effective treatments for such diseases. Recently, automatic diagnosis of Kidney Cancer has become an important challenge especially when using deep learning (DL) due to the importance of training medical datasets, which in most cases are difficult and expensive to obtain. Furthermore, in most cases, algorithms require data from the same domain and a powerful computer with efficient storage capacity. To overcome this issue, a new type of learning known as transfer learning (TL) has been proposed that can produce impressive results based on other different pre-trained data. This paper presents, to the best of the authors’ knowledge, the first comprehensive survey of DL-based TL frameworks for kidney cancer diagnosis. This is a strong contribution to help researchers understand the current challenges and perspectives of this topic. Hence, the main limitations and advantages of each framework are identified and detailed critical analyses are provided. Looking ahead, the article identifies promising directions for future research. Moving on, the discussion is concluded by reflecting on the pivotal role of TL in the development of precision medicine and its effects on clinical practice and research in oncology.
Index Terms:
@@
Artificial intelligence, Deep learning, Domain adaptation, Fine-tuning, Kidney cancer, Transfer learning.
[…]
=-15pt
I Introduction
The kidneys play a crucial role in regulating the body’s fluid balance by filtering out toxins and excreting waste products. Various disorders, such as kidney cancer (KC), can impair kidney function [1]. Fig.1 shows the estimated number of KC incidence and mortality in 2022, for both genders in the world [2]. KC is typically graded on a scale of 1 to 4, with 1 being the least aggressive and 4 being the most aggressive. The grade of a KC is determined by analyzing the cancer cells under a microscope and assessing the degree of differentiation and anaplasia. Grade 1 known as papillary renal cell carcinoma (PRCC), consists of well-differentiated cells closely resembling normal kidney cells [3]. These cancers tend to grow slowly. Grade 2 or chromophobe renal cell carcinoma (CRCC), comprises moderately differentiated cells with some resemblance to normal kidney cells. These cancers grow more slowly than grade 3 or 4 cancers. Grade 3 known as clear cell renal cell carcinoma (CCRCC), consists of poorly differentiated cells bearing little resemblance to normal kidney cells [4]. These cancers tend to grow and spread more rapidly than grade 1 or 2 cancers. Grade 4 called as sarcomatoid renal cell carcinoma (SRCC), is a rare and highly aggressive type of KC that grows and spreads rapidly. It’s important to note that the grade of a KC is one of several factors considered when determining the best course of treatment [5, 6].
AL
artificial learning
@@ I Introduction
DNN
Deep neural network
F1
f1-score
HKS
human kidney segmentation
@@
LLMs
Large language models
@@
UTL
unsupervised TL
CTC
connectionist temporal classification
BERT
bidirectional encoder representations from Transformer
@@
Moreover, KC can be identified by employing a range of imaging methods such as magnetic resonance imaging (MRI), X-rays, and ultrasound (US) imaging [7]. These methods are capable of generating comprehensive visuals of the kidneys, enabling physicians to spot any irregularities or tumors [8]. Nonetheless, despite the utility of medical imagery in disease diagnosis, its reliability is not absolute in certain scenarios. Particularly in cancer cases, more precise techniques may be required to aid physicians in detecting malignant cells and assessing their metastasis to other body regions [9]. Moreover, interpreting whether an image truly signifies the presence of cancer can be challenging. Medical imagery is often rich in detail and data, complicating its analysis. In addition, the overwhelming volume of images and visual assessments may contribute to medical inaccuracies. The burden of analyzing numerous images can diminish a radiologist’s focus and their capacity to notice crucial nuances. Such circumstances can lead to misdiagnosis, diagnosis delays, or even inappropriate treatments [9].
@@
In recent years, deep learning (DL)-based early detection methods have recorded significant achievements in identifying various terminal illnesses [10, 11, 12]. These methods focus on accurately extracting diverse features that assist medical professionals in the diagnostic process through the utilization of the network’s multiple layers [13, 14]. Attaining the necessary diagnostic precision for incurable diseases affecting the kidneys presents challenges due to: (i) the insufficiency of data resources for the training phase [15, 16]; (ii) the time-consuming, expensive, and exhaustive nature of data collection [17, 18]; (iii) the requirement for data annotation in the training phase; and (iv) the impracticality of manually labeling large datasets by experts, which is often unfeasible [19, 20].
@@
While many DL algorithms, such as the convolutional neural network (CNN) [21, 11, 13, 10, 22, 23], are proficient at recognizing complex patterns, they encounter significant challenges, particularly during the transition from training to testing phases, which can impede their effectiveness. Additionally, various studies [24] have pointed out that the majority of datasets utilized for diagnosing either early or advanced cancer stages are often inadequate for the diagnostic process. This inadequacy is due to the intricate structure of certain organs, like the kidneys, and the variations in image depth, making it exceedingly difficult to accurately identify the type or grade of cancer [25, 26, 27]. Furthermore, these algorithms heavily rely on the assumption that the training and testing data are extracted from identical feature spaces and distributions. A deviation between these can necessitate a comprehensive model reconfiguration, which is not only costly but also labor-intensive. As [28] notes, supervised learning algorithms can yield remarkable outcomes with extensive volumes of annotated data. Yet, they generally falter with the complex variations present in medical images, where labeled data are scarce, and backgrounds are complex. Conversely, diagnostic systems powered by artificial intelligence (AI) critically depend on having access to well-distributed and annotated datasets for training, a requirement that is often hard to fulfill [29]. This underscores the need for enhancements in AI-based diagnostic systems for incurable diseases, aiming to simplify them to improve diagnostic efficiency, timing, and precision [30].
@@
To meet these goals, efforts are being directed towards leveraging transfer learning (TL) in many research fields [31, 32, 33, 34, 35]. This process fundamentally involves repurposing a network that has been previously trained on a specific dataset for a distinct task. This entails employing established models such as LeNet, AlexNet, VGGNet, ResNet, GoogLeNet, DenseNet, XceptionNet, and SqueezeNet [32] for tasks on smaller datasets, thereby facilitating the omission of labels. This concept draws inspiration from the human ability to apply knowledge and experiences from past learning to foster new insights. Consequently, various databases hosting extensive image collections have been made available online, such as the CNN algorithm’s application on the ImageNet dataset, which is accessible for public experimentation with DL-based algorithms. This approach effectively mitigates data scarcity issues and circumvents the need for extensive data annotation [36]. Moreover, this algorithmic approach alleviates the lengthy durations typically required for learning processes, which can extend over days or weeks. In the realm of diagnosing terminal illnesses, TL has been categorized into two methodologies: fine-tuning and feature extraction, both of which have shown promising outcomes, notably in KC detection. Numerous studies within this domain, for instance, [37], exemplify its application. TL is particularly beneficial for cancer detection and diagnosis, including KC [38, 39]. Utilizing a pre-trained DL model, like a CNN, fine-tuned on kidney imaging scans (e.g., computerized tomography (CT) or MRI), can enhance its efficacy in identifying and diagnosing kidney tumors (Fig. 2). A method for fine-tuning a model for KC detection might involve using a dataset comprising both healthy and cancerous kidney scans, adjusting the model’s final layers’ weights to better recognize cancerous areas. Additionally, the model could be optimized by either freezing specific layers while training the last few or starting anew with the pre-trained weights. An alternative TL strategy in KC detection could leverage a model pre-tuned for a similar task, such as lung cancer detection, and further fine-tune it on kidney scans, beneficial when large datasets for KC detection are unavailable. TL proves invaluable in medical imaging, significantly when large, labeled datasets are scarce, enhancing the efficiency and accuracy of kidney tumor identification and diagnosis.
Figure 2: Kidney classification based on TL.
I-A Our Contributions
Numerous reviews have covered the contributions of the TL algorithm, citing works such as [37, 40, 41]. However, this paper offers a novel examination by showcasing the role of TL in identifying and managing various types and degrees of renal conditions within a unified and comprehensive framework that encompasses classification and concepts. This approach aids in demystifying the domain of TL and its utilization in KC detection endeavors. The primary focus has been on the enhancement through fine-tuning and domain adaptation (DA) in KC detection, particularly highlighting recent cross-domain advancements. Furthermore, this review underscores the significance of employing TL in data fusion for improving cancer detection and outlines the challenges and future research avenues, suggesting a shift towards generalized TL models for KC detection. The unique contributions of this analysis include:
@@ I-A Our Contributions
  • Exploring various facets of TL and DA that have propelled their evolution.
  • Offering a detailed categorization of TL and DA frameworks along with an exploration of associated concepts.
  • Conducting a thorough literature review on TL and DA’s role in KC detection, grounded in specific standards.
  • Illustrating how TL and DA frameworks enhance integrated data approaches to KC detection.
  • Highlighting unresolved issues in cancer detection faced by TL and DA, such as (1) model precision concerns, (2) selection of evaluation metrics, (3) the challenge of transferring knowledge irrelevant to the subject matter, etc.
  • Proposing future research trajectories for TL and DA frameworks that promise lower computational demands and expenses.
  • Marking this review as a pioneering effort in discussing TL and DA for KC detection, it distinctively contributes by (1) delving into the diverse applications of TL and DA, (2) integrating TL and DA within the KC detection landscape, (3) identifying assorted challenges, and (4) suggesting future research directions.
Table. I, displays the results of the comparison of the contribution of the proposed review with other surveys in the field of TL.
@@
TABLE I: A summary of a comparison of the proposed study’s contribution with other studies in the field of TL is shown. A check mark (✓) signifies that the area has been covered, while a cross mark () indicates that it has not been addressed.
@@
I-B Bibliometric Analysis
@@
To explore and scrutinize the scientific investigations detailed in this review, bibliometric methods were utilized. The cumulative interest in TL-based KC over time is summarized in Fig. 3. This illustration provides a snapshot of the enduring interest in research on TL-based KC, highlighting a notable uptick in enthusiasm for crafting TL-based KC methodologies. Specifically, this surge in interest is represented in Fig. 3(a), indicating a sharp rise in the publication count, which peaked at 66 articles in 2023. Furthermore, Fig.3(b) displays the countries contributing most significantly to research outputs in this area, with China and the United States emerging as the most prolific in TTL-based KC research advancements. Furthermore, Fig. 3(c) identifies the differents domain of application of KC-oriented TL. Lastly, Fig. 3(d) delineates the distribution of paper types, revealing that journal articles predominate, constituting 65.8% of all published work, with conference papers making up 19.7%.
@@
Figure 3: Bibliometric analysis in terms of (a) The number of published articles, (b) The most country known by active researchers, (c) The percentage of the contribution in each domain, and (d) Distribution of paper types.
Table. II, presents a summary of TL research conducted in the field of kidney disease diagnosis.
TABLE II: Summary of TL research conducted in the diagnosis of kidney disease.
Ref. Year Dataset Sample/Input size Used method Performance assessment Objects Methods description KSO
[53] 2022 KiTS - 3D-MS ACC:0.9390 Improve segmentation in Multi-scale approach with a Deep neural network (DNN) Kidney
-RFCNN kidney tumors. based on 3D-MS-RFCNN is applied to capture global
contextual features for kidney tumors
[54] 2022 Private 12,446 VGG16 ACC:99.30 Perform efficient an Development an intelligent system based on EANet, CCT, Kidney
data auto-diagnose kidney diseases Swin transformers and Resnet, VGG16, and Inceptionv3 models
to auto-diagnose kidney stones, cysts, and tumors.
[55] 2022 Private 158 U-Net+TL ${DC}:89.34\pm 5.31$ Improve kidney segmentation Proposed a fully automated 2D attention U-Net model for kidney Kidney
data segmentation on T1 weighted-nephrographic phase contrast enhanced
[56] 2022 Private 162 images CNN+CM DC: 0.853 segmentation of kidney Propose a cross-modal TL from CT to US to Kidney
data +TL in US images segment kidney organ based on CycleGAN, U-Net, U-Res
, PSPNet, and DeepLab v3+ and CNN models.
[57] 2021 Private 258 WSI CNN+TL SEN:0.93 Quantify non-sclerotic and A 9-layer CNN based on the common U-Net Kidney
data F1:0.87 sclerotic glomeruli from kidney. architecture was developed and tested for the segmentation
of non-sclerotic and sclerotic glomeruli of kidney.
[58] 2021 Private - VGG16+TL F1:97.03 Extract the kidney stone The U-Net was used to extract the kidney stone area. Kidney
data fragment area The VGG16 is used as the encoder of U-Net model
to extract the semantic information.
[59] 2020 Private 560/ VGG-16 ACC:0.89 Segmentation of kidneys Investigates the performances of pre-trained U-Net model Kidney
data 768X512 SEN:0.81 from 2D US images. using various backbones for segmentation of
SPE:0.92 kidneys from 2D US images.
[60] 2020 Private 196 FCNN DC:0.88 Segment organs from Used 3D U-Net kernels for segmenting 3D MRI. Kidney
data $VD:0.09\pm 0.05$ large biomedical 3D images
[61] 2019 CHOP 185 kidney TL+CNN ${ACC}:0.989\pm 0.006$ Segment the kidneys Used a boundary distance regression network Kidney
images +BDRN ${DC}:0.94\pm 0.03$ to learn kidney boundary
$MD:2.72\pm 1.61$
[62] 2019 Private 50:CAKUT TL+SVM ACC: 87.00 Classify of normal children support vector machine (SVM) classifiers and TL are used to Kidney
data 50:Controls SEN:86.00 and those with CAKUT. classify diseased and normal kidneys.
[63] 2019 KiTS19 - nnU-Net DC: 0.9620 Kidney segmentation nnU-Net algorithm used to for kidney segmentation Kidney
in CT scans
[64] 2018 Private 48 Slide CNN F1:0.8475 Classified non-sclerosed CNN model is used to classified classifies normal Glomeruli
data images and sclerosed glomeruli. and sclerotic glomeruli
[65] 2018 Private 512x512 Kid-Net F1:Artery:0.72 Segment and classify Propose Kid-Net model for semantic Kidney
data F1:Vein:0.67 kidney vessels kidney segmentation.
[66] 2018 GCDB 935 images DNN+TL ACC:0.8823 Classifying categories of Test DNN+TL model for classifying Glomeruli
SEN:0.8823 glomeruli normal and abnormal categories of glomeruli
[67] 2018 CHOP 50:CAKUT CNN+BDRN $MD:2.8717\pm 2.0665$ Segment the kidneys in Propose subsequent boundary distance regression Kidney
50:CAKUT +PWC ${SEN}:0.9354\pm 0.0625$ clinical US images to segment the kidneys.
[68] 2018 CHOP 50:CAKUT TL+SVM $ACC:0.87\pm 2.1$ Improve CAKUT and kidney Propose a TL-based method and SVM model to extract Kidney
50:Healthy $AUC:0.92\pm 0.7$ diagnosis in children features from US kidney images to improve the CAKUT diagnosis.
Abbreviation: Kind of segmented organ (KSO)
I-C Paper Organization
The structure of this document is outlined below. Section 2 examines the current landscape of TL-driven KC through a detailed classification scheme. Following this, Section 3 outlines the significant advancements in various applications of TL-enhanced KC, such as in the detection of cancer, analysis of CT scans, examination of genomic data, and the integration of multi-modal data. Discussions on the datasets used for KC and the primary benefits and drawbacks of employing TL in the context of KC are presented in Sections 4 and 5, respectively. Subsequently, Section 6 delves into the obstacles faced in KC applications utilizing TL and outlines potential areas for future investigation in Section 7. Conclusions are drawn in the final section, Section 8.
II Taxonomy of TL

### v2

@@ Abstract
Incurable diseases continue to pose significant challenges to global healthcare systems, with their prevalence being influenced by lifestyle choices, economic conditions, social factors, and genetic predispositions. Among these, kidney disease remains a critical health issue, affecting individuals worldwide and necessitating continuous research to enhance early diagnosis and treatment strategies. In recent years, deep learning (DL) has gained prominence in medical imaging and diagnostics, offering significant advancements in automatic kidney cancer (KC) detection. However, the effectiveness of DL models largely depends on the availability of high-quality medical datasets, which are often scarce, expensive, and domain-specific. Additionally, these models require substantial computational resources and storage capacity, limiting their applicability in real-world clinical settings. To address these challenges, transfer learning (TL) has emerged as a promising solution, enabling the adaptation of pre-trained models from different domains to improve KC diagnosis. This paper presents the first comprehensive survey of DL-based TL frameworks specifically applied to KC detection. By systematically reviewing existing approaches, this study identifies key methodologies, explores their advantages and limitations, and critically analyzes their effectiveness in clinical applications. Furthermore, the paper highlights current challenges in implementing TL for medical imaging and discusses emerging trends that could shape future research in this field. Ultimately, this review underscores the transformative role of TL in precision medicine, particularly in oncology, by enhancing diagnostic accuracy, reducing computational costs, and facilitating the deployment of AI-driven healthcare solutions. The findings of this study provide valuable insights for researchers and practitioners, paving the way for further advancements in KC diagnostics and personalized treatment approaches.
keywords:
@@
Medical image, Kidney cancer diagnosis, Deep learning, Transfer learning, Domain adaptation, Fine-tuning.
1 Introduction
The distinction between tumor subtypes and pathological grades is a critical aspect of cancer diagnosis and treatment planning. Tumor subtypes classify the tumor based on its cellular and molecular characteristics. For instance, in renal cell carcinomas (RCC), the major subtypes— clear cell renal cell carcinoma (CCRCC), papillary renal cell carcinoma (PRCC), and chromophobe renal cell carcinoma (CRCC)—differ in their cellular appearance, growth patterns, and genetic alterations. These subtypes have distinct biological behaviors, prognoses, and responses to treatment. RCC is a cancer that impacts kidney function, which plays a vital role in regulating the body’s fluid balance by filtering out toxins and excreting waste products. Disorders such as kidney cancer (KC) can impair this function [1]. As shown in Figure 1, the estimated incidence and mortality rates for KC in 2022, for both genders globally. In terms of pathological grading, KC is classified on a scale of 1 to 4, with 1 being the least aggressive and 4 being the most aggressive. This grading is determined by analyzing cancer cells under a microscope to assess the degree of differentiation and anaplasia. Grade 1 (PRCC) tumors are well-differentiated, closely resembling normal kidney cells and grow slowly [2]. Grade 2 (CRCC) tumors are moderately differentiated with some resemblance to normal kidney cells, growing more slowly than grades 3 and 4. Grade 3 tumors (CCRCC) consist of poorly differentiated cells that bear little resemblance to normal kidney cells and tend to grow and spread rapidly [3]. Grade 4 tumors ( sarcomatoid renal cell carcinoma (SRCC)) are highly aggressive and rare, growing and spreading rapidly. Pathological grading, which reflects how abnormal the tumor cells are compared to normal tissue, helps predict the tumor’s aggressiveness and guides therapeutic decisions. While subtypes categorize the tumor based on its origin and intrinsic characteristics, grading evaluates the degree of malignancy. Understanding both the subtypes and grades is crucial for determining prognosis and selecting the most effective treatment strategy. We have also referenced the WHO pathological classification and grading system for renal tumors [4], providing a reliable and standardized framework for understanding these concepts. These revisions enhance the scientific rigor and clinical relevance of our paper, making it more valuable for researchers and healthcare professionals alike.
Acronyms and Abbreviations
@@ 1 Introduction
BERT
bidirectional encoder representations from Transformer
@@
FSL
few-shot learning
FL
federated learning
@@
LRP
layer-wise relevance propagation
LLM
large language model
@@
UL
unsupervised learning
@@
Moreover, KC can be identified by employing a range of imaging methods such as magnetic resonance imaging (MRI), X-rays, and ultrasound (US) imaging [5]. These methods are capable of generating comprehensive visuals of the kidneys, enabling physicians to spot any irregularities or tumors [6, 7]. Nonetheless, despite the utility of medical imagery in disease diagnosis, its reliability is not absolute in certain scenarios. Particularly in cancer cases, more precise techniques may be required to aid physicians in detecting malignant cells and assessing their metastasis to other body regions [8]. Moreover, interpreting whether an image truly signifies the presence of cancer can be challenging. Medical imagery is often rich in detail and data, complicating its analysis. In addition, the overwhelming volume of images and visual assessments may contribute to medical inaccuracies. The burden of analyzing numerous images can diminish a radiologist’s focus and their capacity to notice crucial nuances. Such circumstances can lead to misdiagnosis, diagnosis delays, or even inappropriate treatments [8].
@@
In recent years, deep learning (DL)-based early detection methods have recorded significant achievements in identifying various terminal illnesses. These methods focus on accurately extracting diverse features that assist medical professionals in the diagnostic process through the utilization of the network’s multiple layers [9, 10]. Attaining the necessary diagnostic precision for incurable diseases affecting the kidneys presents challenges due to: (i) the insufficiency of data resources for the training phase [11, 12]; (ii) the time-consuming, expensive, and exhaustive nature of data collection [13, 14]; (iii) the requirement for data annotation in the training phase; and (iv) the impracticality of manually labeling large datasets by experts, which is often unfeasible [15, 16].
@@
While many DL algorithms, such as the convolutional neural network (CNN) [17, 18, 19, 20], are proficient at recognizing complex patterns, they encounter significant challenges, particularly during the transition from training to testing phases, which can impede their effectiveness. Additionally, various studies [21] have pointed out that the majority of datasets utilized for diagnosing either early or advanced cancer stages are often inadequate for the diagnostic process. This inadequacy is due to the intricate structure of certain organs, like the kidneys, and the variations in image depth, making it exceedingly difficult to accurately identify the type or grade of cancer [22, 23, 24]. Furthermore, these algorithms heavily rely on the assumption that the training and testing data are extracted from identical feature spaces and distributions. A deviation between these can necessitate a comprehensive model reconfiguration, which is not only costly but also labor-intensive. As [25] notes, supervised learning algorithms can yield remarkable outcomes with extensive volumes of annotated data. Yet, they generally falter with the complex variations present in medical images, where labeled data are scarce, and backgrounds are complex. Conversely, diagnostic systems powered by artificial intelligence (AI) critically depend on having access to well-distributed and annotated datasets for training, a requirement that is often hard to fulfill [26]. This underscores the need for enhancements in AI-based diagnostic systems for incurable diseases, aiming to simplify them to improve diagnostic efficiency, timing, and precision [27].
@@
To meet these goals, efforts are being directed towards leveraging transfer learning (TL) in many research fields [28, 29, 30, 31, 32]. This process fundamentally involves repurposing a network that has been previously trained on a specific dataset for a distinct task. This entails employing established models such as LeNet, AlexNet, VGGNet, ResNet, GoogLeNet, DenseNet, XceptionNet, and SqueezeNet [29] for tasks on smaller datasets, thereby facilitating the omission of labels. This concept draws inspiration from the human ability to apply knowledge and experiences from past learning to foster new insights. Consequently, various databases hosting extensive image collections have been made available online, such as the CNN algorithm’s application on the ImageNet dataset, which is accessible for public experimentation with DL-based algorithms. This approach effectively mitigates data scarcity issues and circumvents the need for extensive data annotation [33]. Moreover, this algorithmic approach alleviates the lengthy durations typically required for learning processes, which can extend over days or weeks. In the realm of diagnosing terminal illnesses, TL has been categorized into two methodologies: fine-tuning and feature extraction, both of which have shown promising outcomes, notably in KC detection. Numerous studies within this domain, for instance, [34], exemplify its application. TL is particularly beneficial for cancer detection and diagnosis, including KC [35, 36]. Utilizing a pre-trained DL model, like a CNN, fine-tuned on kidney imaging scans (e.g., computerized tomography (CT) or MRI), can enhance its efficacy in identifying and diagnosing kidney tumors (Figure  2). A method for fine-tuning a model for KC detection might involve using a dataset comprising both healthy and cancerous kidney scans, adjusting the model’s final layers’ weights to better recognize cancerous areas. Additionally, the model could be optimized by either freezing specific layers while training the last few or starting a new one with the pre-trained weights. An alternative TL strategy in KC detection could leverage a model pre-tuned for a similar task, such as lung cancer detection, and further fine-tune it on kidney scans, beneficial when large datasets for KC detection are unavailable. TL proves invaluable in medical imaging, significantly when large, labeled datasets are scarce, enhancing the efficiency and accuracy of kidney tumor identification and diagnosis.
Figure 2: Principle of Kidney classification based on TL.
This study is the first comprehensive review based on DL and TL for the diagnosis of KC. While many studies have explored the application of DL in medical imaging, few have systematically addressed the integration of TL specifically for KC detection. Previous reviews either focus on DL alone or do not consider TL frameworks, which are particularly important for overcoming challenges like dataset scarcity and limited computational resources in real-world clinical settings. This paper fills that gap by presenting a thorough survey of DL-based TL frameworks applied to KC detection, systematically analyzing existing approaches and identifying their advantages and limitations. It goes further by addressing challenges such as the availability of high-quality medical datasets, the need for extensive computational resources, and the difficulties in adapting DL models for practical use in clinical environments. Additionally, the paper highlights emerging trends in TL, which could play a crucial role in enhancing diagnostic accuracy and facilitating AI-driven healthcare solutions. Thus, this review not only addresses a significant gap in the literature but also paves the way for future advancements in KC diagnostics, offering valuable insights for researchers and healthcare practitioners.
1.1 Our Contributions
While numerous review articles have examined the applications of TL algorithms in medical imaging and disease detection—such as [34, 37, 38]—to the best of the authors’ knowledge, no comprehensive survey has been conducted that specifically addresses the role of TL in the context of kidney cancer (KC). This paper aims to fill that gap by offering a novel and unified examination of how TL and domain adaptation ( domain adaptation (DA)) techniques contribute to the detection and management of KC.
Our work provides a structured and in-depth review of TL’s applications across different types and stages of kidney cancer, integrating classification strategies and key conceptual frameworks. This unified perspective aims to demystify the role of TL in KC detection, with a particular focus on enhancements through fine-tuning and cross-domain DA. Furthermore, this review underscores the importance of TL in multi-source data fusion to improve diagnostic accuracy, and it highlights open challenges and future research directions, advocating for the development of more generalizable and computationally efficient TL models for KC detection. The unique contributions of this review are as follows:
@@ 1.1 Our Contributions
  • Exploring various facets of TL and DA that have driven their evolution in medical imaging.
  • Providing a detailed taxonomy of TL and DA frameworks, accompanied by an analysis of their foundational concepts.
  • Conducting a comprehensive literature review on the role of TL and DA in KC detection, based on consistent evaluation criteria. Demonstrating how TL and DA approaches can enhance integrated data pipelines for improved KC diagnosis.
  • Highlighting unresolved challenges in the application of TL and DA to cancer detection, including: (1) model generalizability and precision, (2) evaluation metric selection, (3) the risk of transferring task-irrelevant knowledge, among others.
  • Proposing future research directions aimed at lowering computational complexity and resource demands in TL-based frameworks.
Establishing this work as the first survey, to the best of the authors’ knowledge, to systematically address the integration of TL and DA within the KC detection landscape. It contributes uniquely by (1) investigating diverse applications, (2) presenting an integrated framework, (3) identifying technical and methodological challenges, and (4) offering insights for future exploration. Table. 1, displays the results of the comparison of the contribution of the proposed review with other surveys in the field of TL. This study is intended for a diverse audience that includes researchers, healthcare professionals, and technical experts working in the fields of medical imaging, oncology, and AI. It is particularly relevant for those interested in leveraging cutting-edge DL and TL techniques to improve early detection and diagnosis of KC. This audience may consist of biomedical engineers, data scientists, and oncologists who are looking to enhance diagnostic accuracy, overcome challenges in dataset limitations, and explore innovative methodologies for precision medicine. Furthermore, the study is valuable for academics and industry professionals aiming to integrate advanced AI-driven technologies into clinical practices for cancer diagnosis and treatment planning.
@@
Table 1: A summary of a comparison of the proposed study’s contribution with other studies in the field of TL is shown. A check mark (✓) signifies that the area has been covered, while a cross mark () indicates that it has not been addressed.
@@
2 Methodology
2.1 Search strategy and paper selection
This survey on deep TL for KC diagnosis provides a comprehensive exploration of existing DL-based TL frameworks and their application in the medical field, specifically targeting KC diagnosis. The methodological approach taken in this review encompasses several systematic steps to ensure a thorough analysis and presentation of the current state of research in this area.
The literature search was conducted using several academic databases including PubMed, IEEE Xplore, Scopus, Springer Nature, Elsevier, Wiley, Taylor & Francis, ACM, and Google Scholar. The search was designed to identify relevant articles published in the last five years, ensuring the timeliness and relevance of the data collected. The keywords used for the search included combinations of “TL”, “deep TL”, “DA”, “KC diagnosis”, and “medical imaging”. The search strategy was adapted to the syntax and capabilities of each database to maximize the retrieval of pertinent articles. The main inclusion and exclusion criteria, are:
  • Inclusion criteria:
    • Studies that specifically apply DL and TL techniques.
    • Articles focusing on the diagnosis of KC using medical imaging.
    • Papers published within the last five years to ensure current relevance.
    • Studies providing clear methodological details and quantitative results.
  • Exclusion criteria:
    • Non-English publications.
    • Conference abstracts or presentations without full-text availability.
    • Studies not involving explicit use of DL or TL frameworks.
    • Papers not directly related to KC.
2.2 Bibliometric Analysis
@@
To explore and scrutinize the scientific investigations detailed in this review, bibliometric methods were utilized. The cumulative interest in TL-based KC over time is summarized in Figure 3. This illustration provides a snapshot of the enduring interest in research on TL-based KC, highlighting a notable uptick in enthusiasm for crafting TL-based KC methodologies. Specifically, this surge in interest is represented in Figure 3(a), indicating a sharp rise in the publication count, which peaked at 66 articles in 2023. Furthermore, Figure 3(b) displays the countries contributing most significantly to research outputs in this area, with China and the United States emerging as the most prolific in TTL-based KC research advancements. Furthermore, Figure 3(c) identifies the differents domain of application of KC-oriented TL. Lastly, Figure 3(d) delineates the distribution of paper types, revealing that journal articles predominate, constituting 65.8% of all published work, with conference papers making up 19.7%.
@@
Figure 3: Bibliometric analysis in terms of: (a) The number of published articles; (b) The most country known by active researchers; (c) The percentage of the contribution in each domain; and (d) Distribution of paper types.
Initial screening was based on titles and abstracts, with articles potentially meeting the inclusion criteria selected for full-text review. This stage filtered out irrelevant or out-of-scope studies, streamlining the focus on TL in the context of KC. Essential information was extracted from each article, including study objectives, DL models, datasets, TL techniques employed, and principal findings. This structured approach facilitated comparative analysis and synthesis.
The structure of this document is outlined below. Section 2 describes the methodology used in conducting this review, including the search strategy and bibliometric analysis. Section 3 presents a comprehensive taxonomy of TL and domain adaptation (DA) techniques applied in the context of KC, including inductive, transductive, domain generalization, self-taught, and few-shot learning, along with associated strategies such as fine-tuning and multi-task learning. Section 4 explores key applications of TL in KC, encompassing cancer detection, CT scan analysis and segmentation, genomic data analysis, multi-modal data fusion, and explainable AI. Section 5 provides an overview of the most commonly used datasets and evaluation metrics in this domain. Challenges encountered in employing TL for KC diagnosis and potential directions for future research are discussed in Sections 6 and 7, respectively. Finally, Section 8 summarizes the conclusions drawn from this comprehensive review.
3 Taxonomy of TL